# The influence of the scalar unparticle and polarization on the exclusive W boson hadronic decays in the final state at muon colliders in the Randall-Sundrum model

### v5

@@ Abstract
An attempt is made to present the effect of unparticle physics and polarization on the exclusive decays of W boson at high energy colliders in the Randall-Sundrum (RS) model. By using Feynman diagram techniques we have evaluated the influence of the scalar unparticle and polarization on the exclusive W boson hadronic decays of $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$ and $W^{\pm}\rightarrow\rho^{\pm}\gamma$ at the high energy muon colliders in the RS model. The result shows that with fixed collision energies, the total cross-section for hadronic productions in the final state depends strongly on the parameters of the unparticle physics and muon beam polarizes. With a center-of-mass energy of 10 TeV, the total cross-sections achieve the maximum value when the benchmark signal point as $(\Lambda_{U},d_{U})$ $(1\text{TeV},1.9)$ and the polarization coefficient as $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$. The numerical evaluation for the statistical significance is given in detail which indicates that the effect is greatly enhanced in the future experiments with moderately high energy colliders.
Keywords: scalar unparticle, hadronic decay, muon collider.
@@ I Introduction
     The muon colliders, which can reach center-of-mass energy up to tens of TeV, provide an unprecedented potential in probing new physics beyond the SM[34]. The advantage of initial muon beam polarization is that it is effective for the indirect search [35, 36], which is versatile in the fundamental particle physics and nuclear physics [37]. The integrated luminosity scaling of a high energy muon collider (assuming a 5-year run) reaches 1 $\text{ab}^{-1}$ (3 TeV), 10 $\text{ab}^{-1}$ (10 TeV) and 20 $\text{ab}^{-1}$ (14 TeV)[38, 39, 40, 41]. Given that the radion mixes with the Higgs boson, precision measurements of Higgs properties at muon colliders will provide crucial information [42, 43]. In our previous work, investigation of the scalar unparticle and anomalous couplings at muon colliders in final states with multiple photons in the RS model is considered in detail in Ref.[44]. The contribution of the scalar unparticle on the WW production at ILC is studied in Ref.[45] in which we only considered in s-channel without considering the influence of $\gamma WW,ZWW$ anomalous couplings and neutrino propagator. Moreover, the rare hadronic decays of W bosons in the final states would provide an accurate measurement of the W boson mass that is based solely on visible decay products at future colliders [46, 47, 48]. These rare processes, if observed, would validate the quantum chromodynamics (QCD) factorization formalism used to calculate cross-sections at colliders [48].
     In this work, by using the rare decay channels of $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$, $W^{\pm}\rightarrow\rho^{\pm}\gamma$, we evaluate the influence of the scalar unparticle and polarization on the hadronic production at muon colliders in the RS model. This paper is arranged as follows. The theoretical framework is introduced in Section II. The influence of the scalar unparticle and polarization on the exclusive W boson hadronic decays in the final state at muon colliders are calculated in detail in Section III. Finally, we summarize our results and make conclusions in Section IV.
@@ II Theoretical framework of the Higgs - radion mixing in Randall-Sundrum model
     A compelling question arises regarding whether a intuitive and tractable alternative exists for modeling unparticles. A natural theoretical candidate for this purpose is found in models based on warped extra dimensions. This idea is mentioned in RS model [28, 29]. We use couplings of the SM particles with the Higgs, radion and unparticle listed in detail in Ref.[5, 19, 17, 14] for evaluating in the next section.
III The influence of the scalar unparticle and polarization on the exclusive W boson hadronic decays in the final state at muon colliders
@@
     Study for the exclusive boson hadronic decay with the ATLAS detector at 13 TeV is considered in detail in Ref.[48]. Recently, mass biases in reconstructing exclusive radiative hadronic decays of W boson at the LHC are given in Ref.[50]. While rare and exclusive few-body decays of the Higgs, Z and W bosons and the top quark are studied in Ref.[51], the W boson’s exotic decay channel is concerned in Ref.[52]. However, the influence of the unparticle and polarization at muon colliders through hadronic decays of W boson has not yet been invested. In this section, by using the rare decay channels of $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$, $W^{\pm}\rightarrow\rho^{\pm}\gamma$, we investigate the influence of the scalar unparticle and polarization on the hardronic production at muon colliders in the RS model.
@@
where $\psi=(\overrightarrow{p}_{1},\overrightarrow{k}_{1})$ is the scattering angle. The model parameters are chosen as $\lambda_{\mu\mu}=\lambda_{WW}=\lambda_{0}=1$, $\xi=1/6$ [5], $\Lambda_{\phi}=5$ TeV [6], $m_{h}$ = 125 GeV, $m_{\phi}$ = 125 GeV [33]. The integrated luminosity scaling of a high energy muon collider reaches 10 $\text{ab}^{-1}$ at a center-of-mass energy of $\sqrt{s}=10$ TeV [40, 41]. The bounds on the anomalous $W^{-}W^{+}\gamma$ and $W^{-}W^{+}Z$ couplings are provided by the LEP, Tevatron and LHC experiments. The ATLAS collaboration has updated the best available constraints on anomalous couplings $\Delta k_{\gamma}$, $\lambda_{\gamma}$, $\Delta k_{Z}$, $\lambda_{Z}$ obtained as follows: $\Delta k_{\gamma}\in[-0.135,0.190]$, $\lambda_{\gamma}\in[-0.065,0.061]$, $\Delta k_{Z}\in[-0.061,0.093]$, $\lambda_{Z}\in[-0.062,0.065]$ [56]. From the formulas (28 - 30) above and branching ratio of W boson in the soft-collinear effective theory (SCET) and the light-cone distribution amplitude (LCDA) framework shown in Ref.[51], we give estimates in detail for the cross-sections as follows:
    i) We evaluate the benchmark parameters $(\Lambda_{U},d_{U})$ in Fig.2. The parameters are taken to be $P_{\mu^{-}}=0.8,P_{\mu^{+}}=-0.8$, $\sqrt{s}=10$ TeV, $\Delta k_{\gamma}=0.190$, $\lambda_{\gamma}=0.061$, $\Delta k_{Z}=-0.061$, $\lambda_{Z}=-0.062$. From the figure, we can see that the cross-sections reach their maximum values at the benchmark signal point $(\Lambda_{U},d_{U})$ $(1\text{TeV},1.9)$. This result is similar to Ref.[32].
@@
ii) With the benchmark signal point $(\Lambda_{U},d_{U})=(1\text{TeV},1.9)$, the total cross-section depends on ($\Delta k_{\gamma}$, $\lambda_{\gamma}$) shown in the Fig. 3. The other parameters are chosen as in Fig.2. The figure shows that cross-sections are the largest in the yellow region, the typical values are given by (a) $0.8\times 10^{-9}$ fb for the decay $W^{\pm}\rightarrow\pi^{\pm}\gamma$, (b) $0.53\times 10^{-11}$ fb for the decay $W^{\pm}\rightarrow K^{\pm}\gamma$ and (c) $3.8\times 10^{-9}$ fb for the decay $W^{\pm}\rightarrow\rho^{\pm}\gamma$, respectively.
@@
iii) Similar to the Fig.3, the total cross-sections depend on ($\Delta k_{Z}$, $\lambda_{Z}$) are shown in the Fig.4. The typical values of the cross -section are given by (a) $0.810\times 10^{-9}$ fb, (b) $0.535\times 10^{-11}$ fb and (c) $3.870\times 10^{-9}$ fb, respectively.
@@
iv) In Fig.5, the cross-sections are plotted as the function of polarization coefficients ($P_{\mu^{-}},P_{\mu^{+}}$). The parameters are chosen as above, i.e $d_{U}=1.9$, $\Lambda_{U}=1$ TeV, $\sqrt{s}=10$ TeV, $\Delta k_{\gamma}=0.190$, $\lambda_{\gamma}=0.061$, $\Delta k_{Z}=-0.061$, $\lambda_{Z}=-0.062$. The figures indicate that the cross-sections achieve the maximum values when $P_{\mu^{-}}=P_{\mu^{+}}=1$ or 1-1 and the minimum values when $P_{\mu^{-}}=1,P_{\mu^{+}}=-1$ or $P_{\mu^{-}}=-1,P_{\mu^{+}}=1$, respectively.
@@
v) The cross-section dependence on the energy is shown in Fig.6. The parameters are chosen as in Fig.5 and the polarization coefficients $(P_{\mu^{-}},P_{\mu^{+}})$ are similar to Ref.[49], i.e $(P_{\mu^{-}},P_{\mu^{+}})$ =(1,1),(0.8,0.8),(0.6,0.6),(1,1)=(1,-1),(0.8,-0.8),(0.6,-0.6),(1,1), respectively. From the figure we can see that the cross-sections increase when the collision energy increases. In case of $\mu^{-}$ beam is left polarized, $\mu^{+}$ beam is right polarized and vice versa, the cross-sections change insignificantly.
@@
vi) Finally, for more details, we calculate the statistical significance for the exclusive W boson hadronic decays at muon colliders which is defined by [57, 58] $SS=S/\sqrt{S+B}$, where B, S are the number of events for the signal in SM framework and in RS framework including unparticle, anomalous couplings, respectively. The parameters are chosen as in Fig.5 and the polarization coefficients of $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$ are chosen. Using the integrated luminosity scaling of 10 $\text{ab}^{-1}$ (assuming a 5-year run) [40, 41], the some typical values of the statistical significance are given in Table.1. The analysis reveals that the statistical significance of the exclusive hadronic W boson channel at muon colliders remains small, yielding $\text{SS}=(0.0048\pm 0.00096)\sigma$, $(0.00039\pm 0.00008)\sigma$, and $(0.0105\pm 0.00229)\sigma$ for the decays of $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$ and $W^{\pm}\rightarrow\rho^{\pm}\gamma$, respectively. However, with the contribution from the RS model including unparticle, and anomalous couplings, the cross-sections are greatly enhanced (the readers can see more details in Table.2). It is worth that there is not universal agreement between these predictions. For example, the statistical significance SS can reach $3\sigma-8\sigma$ in case of using the W branching ratio at the experimental upper bounds of $\mathcal{O}(10^{-6})$ in Ref.[48].
Table 1: Some typical values of the statistical significance for the exclusive W boson hadronic decays at high energy muon colliders with the integrated luminosity of 10 $\text{ab}^{-1}$. The parameters are chosen as in Fig.5
and the polarization coefficients of $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$ are chosen.
@@
Table 2: Some typical values in the layers: SM, RS including unparticle and anomalous couplings. The parameters are taken to be $(\Lambda_{U},d_{U})$ $(1\text{TeV},1.9)$ and the polarization coefficients of $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$ are chosen.

### v6

# The contribution of new physics on the exclusive W boson hadronic decays in the final state at muon colliders in the Randall-Sundrum model

@@ Abstract
An attempt is made to present the effect of new physics in the exclusive decays of W boson at high energy colliders in the Randall-Sundrum (RS) model. By using Feynman diagram techniques we have evaluated the influence of the scalar unparticle, polarization and anomalous couplings on the exclusive W boson hadronic decays of $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$ and $W^{\pm}\rightarrow\rho^{\pm}\gamma$ at the high energy muon colliders in the RS model. The result shows that with fixed collision energies, the total cross-section for hadronic productions in the final state depends strongly on the parameters of the unparticle physics, muon beam polarizes and also anomalous couplings. With a center-of-mass energy of 10 TeV, the total cross-sections achieve the maximum value when the benchmark signal point as $(\Lambda_{U},d_{U})$ $(1\text{TeV},1.9)$ and the polarization coefficient as $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$. The numerical evaluation shows that with the contribution of new physics in the RS model, the effect is greatly enhanced at high energy colliders. The statistical significance can reach $1\sigma-7\sigma$ in case of using the W branching ratio at the experimental bounds of $\mathcal{O}(10^{-6})$. To clarify the contribution of new physics, we use a $\chi^{2}$ analysis with systematic errors to determine the sensitivities of the new contributions. The result indicates that the sensitivities on the anomalous coupling $WW\gamma$ are much larger than that on the anomalous coupling $WWZ$ under the same conditions.
Keywords: scalar unparticle, anomalous coupling, hadronic decay, muon collider.
@@ I Introduction
     The muon colliders, which can reach center-of-mass energy up to tens of TeV, provide an unprecedented potential in probing new physics beyond the SM[34]. The advantage of initial muon beam polarization is that it is effective for the indirect search [35, 36], which is versatile in the fundamental particle physics and nuclear physics [37]. The integrated luminosity scaling of a high energy muon collider (assuming a 5-year run) reaches 1 $\text{ab}^{-1}$ (3 TeV), 10 $\text{ab}^{-1}$ (10 TeV) and 20 $\text{ab}^{-1}$ (14 TeV)[38, 39, 40, 41]. Given that the radion mixes with the Higgs boson, precision measurements of Higgs properties at muon colliders will provide crucial information [42, 43]. In our previous work, investigation of the scalar unparticle and anomalous couplings at muon colliders in final states with multiple photons in the RS model is considered in detail in Ref.[44]. The contribution of the scalar unparticle on the WW production at ILC is studied in Ref.[45] in which we only considered in s-channel without considering the influence of $\gamma WW,ZWW$ anomalous couplings and neutrino propagator. Moreover, the rare hadronic decays of W bosons in the final states would provide an accurate measurement of the W boson mass that is based solely on visible decay products at future colliders [46, 47, 48]. These rare processes, if observed, would validate the quantum chromodynamics (QCD) factorization formalism used to calculate cross-sections at colliders [48]. In our work, we consider the exclusive decays of W boson into a photon plus a charged meson. It is worth that there is not universal agreement between these predictions. The theoretical branching fractions are in the $\mathcal{O}(10^{-8}-10^{-13})$ range [49, 50], while the current experimental bounds are larger in the range of $\mathcal{O}(10^{-4}-10^{-6})$ [47, 48, 50].
     In this work, by using the rare decay channels of $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$, $W^{\pm}\rightarrow\rho^{\pm}\gamma$, we evaluate the contribution of new physics on the hadronic production at muon colliders in the RS model. This paper is arranged as follows. The theoretical framework is introduced in Section II. The contribution of new physics on the exclusive W boson hadronic decays in the final state at muon colliders are calculated in detail in Section III. Finally, we summarize our results and make conclusions in Section IV.
@@ II Theoretical framework of the Higgs - radion mixing in Randall-Sundrum model
     A compelling question arises regarding whether a intuitive and tractable alternative exists for modeling unparticles. A natural theoretical candidate for this purpose is found in models based on warped extra dimensions. This idea is mentioned in RS model [28, 29]. The effective Lagrangian can be written by
$\mathcal{L}_{eff}=\mathcal{L}_{SM}+\mathcal{L}_{RS}(h,\phi;\xi,\Lambda_{\phi})+\frac{C_{UW}}{\Lambda_{U}^{d_{U}}}\mathcal{O}_{U}\mathcal{O}_{W}+\frac{C_{U\mu}}{\Lambda_{U}^{d_{U}-1}}\mathcal{O}_{U}\overline{\mu}\mu+$... (16)
We use couplings of the SM particles with the Higgs, radion and unparticle listed in detail in Ref.[5, 19, 17, 14] for evaluating in the next section.
III The contribution of new physics on the exclusive W boson hadronic decays in the final state at muon colliders
@@
     Study for the exclusive boson hadronic decay with the ATLAS detector at 13 TeV is considered in detail in Ref.[48]. Recently, mass biases in reconstructing exclusive radiative hadronic decays of W boson at the LHC are given in Ref.[52]. While rare and exclusive few-body decays of the Higgs, Z and W bosons and the top quark are studied in Ref.[50], the W boson’s exotic decay channel is concerned in Ref.[53]. However, the contribution of unparticle and anomalous couplings at muon colliders through hadronic decays of W boson has not yet been invested. In this section, by using the rare decay channels of $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$, $W^{\pm}\rightarrow\rho^{\pm}\gamma$, we investigate the influence of new physics on the hadronic production at muon colliders in the RS model.
@@
where $\psi=(\overrightarrow{p}_{1},\overrightarrow{k}_{1})$ is the scattering angle. The model parameters are chosen as $\lambda_{\mu\mu}=\lambda_{WW}=\lambda_{0}=1$, $\xi=1/6$ [5], $\Lambda_{\phi}=5$ TeV [6], $m_{h}$ = 125 GeV, $m_{\phi}$ = 125 GeV [33]. The integrated luminosity scaling of a high energy muon collider reaches 10 $\text{ab}^{-1}$ at a center-of-mass energy of $\sqrt{s}=10$ TeV [40, 41]. The bounds on the anomalous $W^{-}W^{+}\gamma$ and $W^{-}W^{+}Z$ couplings are provided by the LEP, Tevatron and LHC experiments. The ATLAS collaboration has updated the best available constraints on anomalous couplings $\Delta k_{\gamma}$, $\lambda_{\gamma}$, $\Delta k_{Z}$, $\lambda_{Z}$ obtained as follows: $\Delta k_{\gamma}\in[-0.135,0.190]$, $\lambda_{\gamma}\in[-0.065,0.061]$, $\Delta k_{Z}\in[-0.061,0.093]$, $\lambda_{Z}\in[-0.062,0.065]$ [57]. From the formulas (28 - 30) above and branching ratio of W boson in the soft-collinear effective theory (SCET) and the light-cone distribution amplitude (LCDA) framework shown in Ref.[50], we give estimates in detail for the cross-sections from Fig.2 to Fig.6 as follows:
    Initially, we evaluate the benchmark parameters $(\Lambda_{U},d_{U})$ in Fig.2. The parameters are taken to be $P_{\mu^{-}}=0.8,P_{\mu^{+}}=-0.8$, $\sqrt{s}=10$ TeV, $\Delta k_{\gamma}=0.190$, $\lambda_{\gamma}=0.061$, $\Delta k_{Z}=-0.061$, $\lambda_{Z}=-0.062$. From the figure, we can see that the cross-sections reach their maximum values at the benchmark signal point $(\Lambda_{U},d_{U})$ $(1\text{TeV},1.9)$.
@@
Furthermore, with the benchmark signal point $(\Lambda_{U},d_{U})=(1\text{TeV},1.9)$, the total cross-section depends on ($\Delta k_{\gamma}$, $\lambda_{\gamma}$) shown in the Fig. 3. The other parameters are chosen as in Fig.2. The figure shows that cross-sections are the largest in the yellow region, the typical values are given by (a) $0.8\times 10^{-9}$ fb for the decay $W^{\pm}\rightarrow\pi^{\pm}\gamma$, (b) $0.53\times 10^{-11}$ fb for the decay $W^{\pm}\rightarrow K^{\pm}\gamma$ and (c) $3.8\times 10^{-9}$ fb for the decay $W^{\pm}\rightarrow\rho^{\pm}\gamma$, respectively.
@@
Similar to the Fig.3, the total cross-sections depend on ($\Delta k_{Z}$, $\lambda_{Z}$) are shown in the Fig.4. The typical values of the cross -section are given by (a) $0.810\times 10^{-9}$ fb, (b) $0.535\times 10^{-11}$ fb and (c) $3.870\times 10^{-9}$ fb, respectively.
@@
In Fig.5, the cross-sections are plotted as the function of polarization coefficients ($P_{\mu^{-}},P_{\mu^{+}}$). The parameters are chosen as above, i.e $d_{U}=1.9$, $\Lambda_{U}=1$ TeV, $\sqrt{s}=10$ TeV, $\Delta k_{\gamma}=0.190$, $\lambda_{\gamma}=0.061$, $\Delta k_{Z}=-0.061$, $\lambda_{Z}=-0.062$. The figures indicate that the cross-sections achieve the maximum values when $P_{\mu^{-}}=P_{\mu^{+}}=1$ or 1-1 and the minimum values when $P_{\mu^{-}}=1,P_{\mu^{+}}=-1$ or $P_{\mu^{-}}=-1,P_{\mu^{+}}=1$, respectively.
@@
Next, we evaluate the dependence of the cross-section on the energy, as shown in Fig.6. The parameters are chosen as in Fig.5 and the polarization coefficients $(P_{\mu^{-}},P_{\mu^{+}})$ are similar to Ref.[51], i.e $(P_{\mu^{-}},P_{\mu^{+}})$ =(1,1),(0.8,0.8),(0.6,0.6),(1,1)=(1,-1),(0.8,-0.8),(0.6,-0.6),(1,1), respectively. From the figure we can see that the cross-sections increase when the collision energy increases. In case of $\mu^{-}$ beam is left polarized, $\mu^{+}$ beam is right polarized and vice versa, the cross-sections change insignificantly.
@@
For more details, we calculate the statistical significance for the exclusive W boson hadronic decays at muon colliders which is defined by [58, 59] $SS=S/\sqrt{S+B}$, where B, S are the number of events for the signal in SM framework and in RS framework including unparticle, anomalous couplings, respectively. The parameters are chosen as in Fig.5 and the polarization coefficients of $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$ are chosen. Using the integrated luminosity scaling of 10 $\text{ab}^{-1}$ (assuming a 5-year run) [40, 41], the some typical values of the statistical significance with different values of the W branching ratio in SCET and LCDA framework in Ref.[50] are given in detail in Tables 1 and in Table 2, respectively. From Table 1 we can see that at the theoretical bound of $\mathcal{O}(10^{-9})$, the statistical significance remains small. While in Table 2, using the W branching ratio at the experimental bounds of $\mathcal{O}(10^{-6})$[48] the statistical significance is given by, $\text{SS}=(1.4444^{+0.6018}_{-0.3611})\sigma$, $(1.3239^{+0.4815}_{-0.3611})\sigma$ and $(7.2218^{+2.7684}_{-2.0463})\sigma$ for the decays of $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$ and $W^{\pm}\rightarrow\rho^{\pm}\gamma$, respectively. This result shows that with the contribution the new physics from the RS model including unparticle and anomalous couplings the effect is enhanced and can be measured at the multi-TeV muon collider.
Table 1: Some typical values of the statistical significance for the exclusive W boson hadronic decays at high energy muon colliders with the integrated luminosity of 10 $\text{ab}^{-1}$ in case of using the W branching ratio in the SCET and LCDA framework in Ref.[50], $\mathcal{O}(10^{-9})$. The parameters are chosen as in Fig.5 and the polarization coefficients of $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$ are chosen.
@@
Table 2: Some typical values in the layers: SM, RS including unparticle and anomalous couplings in case of using the W branching ratio in the SCET and LCDA framework in Ref.[50], $\mathcal{O}(10^{-9})$. The parameters are chosen as in Fig.5 and the polarization coefficients of $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$ are chosen.
@@
Table 3: Some typical values of the statistical significance for the exclusive W boson hadronic decays at high energy muon colliders with the integrated luminosity of 10 $\text{ab}^{-1}$ in case of using the W branching ratio at the experimental bounds of $\mathcal{O}(10^{-6})$ in Ref.[48]. The parameters are chosen as in Fig.5 and the polarization coefficients of $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$ are chosen.
Process $\sigma_{B}$ (fb) $\sigma_{S}$ (fb) $SS(\sigma)$
$\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\pi^{-}\pi^{+}\gamma\gamma$ ($3.4407^{+0.5973}_{-0.2150}$ ) $\times 10^{-4}$ (${3.9182}^{+0.6802}_{-0.2449}$) $\times 10^{-4}$ $1.4444^{+0.6018}_{-0.3611}$
$\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow K^{-}K^{+}\gamma\gamma$ ($2.8911^{+0.3823}_{-0.2150}$) $\times 10^{-4}$ ($3.2923^{+0.4354}_{-0.2449}$) $\times 10^{-4}$ $1.3239^{+0.4815}_{-0.3611}$
$\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\rho^{-}\rho^{+}\gamma\gamma$ ($8.6016^{+1.2639}_{-0.6905}$) $\times 10^{-3}$ ($9.7954^{+1.4394}_{-0.7864}$) $\times 10^{-3}$ $7.2218^{+2.7684}_{-2.0463}$
Finally, to clarify the contribution of the anomalous couplings $WW(\gamma,Z)$, we use a $\chi^{2}$ analysis with systematic errors to determine the sensitivities of the new contributions at the multi-TeV muon colliders. The $\chi^{2}$ function is defined as follows [61, 60]
$\chi^{2}=\left(\dfrac{\sigma_{SM}-\sigma_{NP}}{\sigma_{SM}\sqrt{(\delta_{st})^{2}+(\delta_{sys})^{2}}}\right)^{2},$ (32)
where $\sigma_{SM}$ is the cross-section of SM backgrounds, $\sigma_{NP}$ is the cross-section containing contributions from presence of both new physics beyond the SM and SM backgrounds, defined by $\sigma_{NP}=\sigma_{SM}+\sigma_{INT}+\sigma_{BSM}$. Here, $\sigma_{INT}$ is the interference term between the SM and the new physics contribution and $\sigma_{BSM}$ is the contribution due to BSM physics, respectively. $\delta_{st}=\dfrac{1}{\sqrt{N_{SM}}}=\dfrac{1}{\sigma_{SM}\times\mathcal{L}}$ and $\delta_{sys}$ are the statistical error and the systematic error, respectively. $N_{SM}=\sigma_{SM}\times\mathcal{L}$ is the number of events of SM backgrounds, $\mathcal{L}$ is the integrated luminosity. The parameters are chosen as in Table.2. The systematic error is chosen as $\delta_{sys}=5\%$ [61]. Some typical values of $\chi^{2}$ for the contributions of the anomalous coupling $WW(\gamma,Z)$ are evaluated in case of using the W branching ratio of $\mathcal{O}(10^{-6})$ in Table.4. The sensitivities to scalar propagators (unparticle, radion and Higgs) are found to be negligible. The best sensitivities obtained from the processes $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\pi^{-}\pi^{+}\gamma\gamma/K^{-}K^{+}\gamma\gamma/\rho^{-}\rho^{+}\gamma\gamma$ on the anomalous coupling $WW\gamma$ are much larger than that of the anomalous coupling $WWZ$ under the same conditions.
Table 4: Some typical values of $\chi^{2}$ for the contributions of photon, Z boson propagators in case of using the W branching ratio at the experimental upper bounds of $\mathcal{O}(10^{-6})$ in Ref.[48]. The parameters are chosen as in Table.2. The systematic error is chosen as $\delta_{sys}=5\%$.
Final states $\sqrt{s}$ (TeV) $\mathcal{L}$ ($ab^{-1}$) $\chi_{\gamma}^{2}$ $\chi_{Z}^{2}$
$\pi^{-}\pi^{+}\gamma\gamma$ 3 1 2.3641 $\times 10^{-4}$ 5.9326 $\times 10^{-7}$
10 10 2.5246 $\times 10^{-2}$ 6.1945 $\times 10^{-5}$
14 20 9.3737 $\times 10^{-2}$ 2.2975 $\times 10^{-4}$
$K^{-}K^{+}\gamma\gamma$ 3 1 1.8406 $\times 10^{-4}$ 4.6189 $\times 10^{-7}$
10 10 1.9729 $\times 10^{-2}$ 4.8409 $\times 10^{-5}$
14 20 7.4022 $\times 10^{-2}$ 1.8143 $\times 10^{-4}$
$\rho^{-}\rho^{+}\gamma\gamma$ 3 1 5.6155 $\times 10^{-3}$ 1.4092 $\times 10^{-5}$
10 10 4.3371 $\times 10^{-1}$ 1.0642 $\times 10^{-3}$
14 20 9.1075 $\times 10^{-1}$ 2.2322 $\times 10^{-3}$